OpenStudy (anonymous):

Your valentine gives you a box of chocolate covered candy for Valentine's Day. The box contains 10 square pieces of candy; 1 strawberry, 2 coconut, 3 orange, and 4 chocolate. What is the probability that you eat a chocolate candy and then a strawberry candy? 2.Scott built a pond and stock it with 130 bass, 75 crappie, 210 bream, 30 catfish, and 55 shell cracker. After he released the fish into the pond, he decided to go fishing. He decided that he would return any fish that he caught to the pond (catch and release). What is the probability that he caught two bass and a bream?

OpenStudy (anonymous):

Hey thanks for helping lucaz

OpenStudy (lucaz):

1. you have a total of 10 pieces of candy, P(chocolate and strawberry) is: P(c) * P(s) = (4/10) * (1/9) = 2/45 2.P(2 bass and bream)= P(bass) * P(bass) * P(bream)= (130/500) * (130/500) * (210/500) since he decided to return the fish the total keeps the same.

OpenStudy (anonymous):

thanks

OpenStudy (lucaz):

the idea is the (# of what you want) / (# of all possibilities)

OpenStudy (anonymous):

ok got it i made a 100 thanks so much

OpenStudy (lucaz):

you're welcome

OpenStudy (anonymous):

Do you mind helping me with some more like this?